OpenStudy (anonymous):

simplify 5 sqrt v^8 - sqrt v^3 - sqrt 16v^8 + 3 sqrt v^3

OpenStudy (cwrw238):

evalute like terms eg - sqrt v^3 + 3 sqrt v^3 = ?

OpenStudy (cwrw238):

also sqrt 16v^8 = 4 sqrt v^8

OpenStudy (anonymous):

i thought you had to simplify first ?

OpenStudy (anonymous):

\[5\sqrt{v^{8}} - \sqrt{v^{3}} - 4\sqrt{v^{8}} + 3\sqrt{v^{3}}\]

OpenStudy (anonymous):

That becomes what cwrw was saying.

OpenStudy (cwrw238):

thats what i'm doing what is 5 sqrt v^8 - 4 sqrt v^8 ?

OpenStudy (anonymous):

1 sqrt v^8?

OpenStudy (anonymous):

1 sqrt v^8?

OpenStudy (cwrw238):

yep now work out whats in my first post

OpenStudy (cwrw238):

- sqrt v^3 + 3 sqrt v^3 =

OpenStudy (anonymous):

2 sqrt v^3?

OpenStudy (cwrw238):

right now put them together and you have it

OpenStudy (cwrw238):

one term iv v^3 and one term in v^8

OpenStudy (cwrw238):

the answer is 2 sqrt v^3 + sqrt v^8

OpenStudy (anonymous):

thats not one of my options ): @cwrw238

OpenStudy (cwrw238):

ok what are they?

OpenStudy (anonymous):

v^4+2v sqrt v, 9v^4 sqrt v^3, 9v^6 sqrt v, and 2v sqrt v^5 @cwrw238

OpenStudy (cwrw238):

ok - sorry I should have simplified further ( my bad decision0 v^8 = v^4 * v^4 so sqrt v^8 = v^4

OpenStudy (cwrw238):

2 sqrt v^3 = 2* sqrt v^2 * sqrt v = 2*v * sqrt v answer is v^4 + 2v sqrt v

OpenStudy (anonymous):

thank you (:

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

sometimes we can be a bit too confident.....